Notyourfreind:

What is f(5) if f(1) = 3.2 and f(x + 1) = Five-halves(f(x))?

mdobbs6856:

Can you explain the problem?

Notyourfreind:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @mdobbs6856 Can you explain the problem? \(\color{#0cbb34}{\text{End of Quote}}\) That's all it tells me... and I'm bad at math

mdobbs6856:

then f(2) = f(2-1) - 5 = f(1) - 5 = 2 - 5 = -3 thenf(3) = f(3-1) - 5 = f(2) - 5 = -3 - 5 = -8 then f(4) = f(4-1) - 5 = f(3) - 5 = -8 - 5 = -13 then f(5) = f(5-1) - 5 = f(4) - 5 = -13 -5 = -18 then f(6) = f(6-1) - 5 = f(5) - 5 = -18 -5 = -23